OpenStudy (anonymous):

I have 2 sides of a irregular tria angle. I have the angle degrees. What formula do I use to calculate the length of the third side.

OpenStudy (amistre64):

law of cosines:

OpenStudy (amistre64):

if the angle is tucked between the legs like a whimpering dog, use law of cosines

OpenStudy (amistre64):

c^2 = a^2 + b^2 -2ab cos(C)

OpenStudy (amistre64):

\[c =\sqrt{a^2 + b^2 -2ab \cos(C)}\]